evaluate ∫ (1 – cos 2x) / (1 + cos 2x) dx

evaluate
∫  (1 – cos 2x) / (1 + cos 2x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

I = \int \frac{1-cos2x}{1+cos2x}dx
I = \int \frac{2sin^{2}x}{2cos^{2}x}dx
I = \int \frac{sin^{2}x}{cos^{2}x}dx
I = \int tan^{2}xdx
I = \int (sec^{2}x-1)dx
I = tanx + c